OpenStudy (anonymous):

The time to complete a project is inversely proportional to number of people who are doing the project. A certain project can be completed by 5 workers in 24 days. In order to finish the project sooner, the company plans to hire additional workers. How many workers are needed to finish the project in 15 days?

OpenStudy (anonymous):

Write an inverse variation equation and solve the problem.

OpenStudy (anonymous):

Let's solve for d. 5w+24d=15 Step 1: Add -5w to both sides. 24d+5w+−5w=15+−5w 24d=−5w+15 Step 2: Divide both sides by 24. 24d24=−5w+1524 d=−524w+58 Answer: d=−524w+58

OpenStudy (anonymous):

but ai'nt you gottah solve that

OpenStudy (anonymous):

oh my fought nvm i got it thanks aloud